[Brahms-l] Shift report 20050211 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Feb 11 2005 - 00:02:40 EST
1. To stores. First from 16.45 - 21.25. Last from 23.00 - scheduled for dump at 03.50.
2 MRS@45deg@A1050
FS@3deg@B3450
run      tr2       tr3
13067    224k      123k
13068    137k      123k
13069    288k      218k
13070     84k       99k
Non-TPC pedestal - run 13071
TPC pedestal - run 13084
run 13085 in progress
3. Had same problem when I tried to do TPC pedestal, but Hiro fixed it.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Feb 11 00:03:11 2005

This archive was generated by hypermail 2.1.8 : Fri Feb 11 2005 - 00:03:25 EST